# SMSupplementary References

### v1

@@ 5.3 Is a good classifier all you need for VLM Open-set Recognition?
Classification: The correlation observed by [43] appears to hold for VLM classifiers, with greater accuracy (closed-set performance) relating to greater AuPR scores (open-set performance). This trend is also confirmed in Fig. 6, which compares closed-set and open-set performance given different query set sizes. It is worth noting that AuPR is a summary metric and this behaviour can be inconsistent when examining performance at different operating points on the precision-recall curves. For example, ImageBind shows the greatest accuracy and AuPR, but it is outperformed by ALIGN by $3\%$ recall at the $95\%$ precision point.
@@ 5.4 How Does Query Set Size Impact Performance?
Figure 6: Influence of the query set size on closed-set and open-set performance in VLM image classification. Random subsets of ImageNet1k classes were selected for the query set, with mean and standard deviation from 10 random seeds visualised.
In Fig. 6, we test the VLM classifiers with different query set sizes by randomly sampling subsets of target classes from the ImageNet dataset. Increasing the number of target classes and thus increasing the query set size consistently decrements both the closed-set and open-set performance of a VLM classifier. This indicates that testing VLMs with an entire vocabulary of nouns may not be a feasible approach to avoid open-set recognition.
6 Discussion and Conclusions
The goal of this work is to highlight the vulnerability of VLMs for open-set recognition and enable future research in this direction. In this spirit, we now highlight the research directions that emerge from our results for future work.
Uncertainty Measures for Identifying Open-set Error: The baseline uncertainty measures from VLMs (softmax/sigmoid, cosine, and entropy) were not effective for distinguishing between true positive predictions and open-set errors. This indicates a need for alternative uncertainty measures well-suited to this task. While there have been some works exploring uncertainty representations for VLMs [20, 42], their utility for open-set recognition remains unexplored.
Selecting Optimal Negative Embeddings: Our random placement of negative embeddings required large numbers of negative embeddings to induce any effect and exhibited a trade-off where negative embeddings also captured true positive predictions. After additional analysis, we observe that some negative embeddings are more effective than others at capturing open-set errors – for example, testing with ImageBind, the random string “awy” captured ten times more open-set errors than true positive predictions (see Supplementary Material). Identifying the characteristics of these embeddings or learning them through prompt-tuning [14, 12, 9, 58, 23] is an interesting direction.
Application-specific Open-set Recognition for VLMs: Our work tests on popular, general-purpose computer vision datasets. Application-specific datasets can be more realistic for the actual deployment of computer vision models and can contain conditions that may not be represented in general-purpose datasets, e.g. poor visual conditions, semantic overlap between classes, class imbalance, etc. How these factors influence the open-set problem for VLMs is an important direction that warrants further investigation by the community.
Open-set Error Cost: A limitation of our work is that we treat every open-set error equally and do not consider the semantic distance between confused classes (i.e. mistaking a car as a truck is as bad as misclassifying a dog as a truck). Although a more nuanced error measure might be beneficial for some applications, it is beyond the scope of our paper, but interesting for future work.
@@ 6 Discussion and Conclusions
[…]
@@
The area under the ROC curve (AuROC) is a popular metric in the traditional open-set recognition literature [8, 30, 32, 43, 54]. However a key limitation of ROC curves is that they are not affected by the size of the negative class [39] – this can make them less informative and even misleading when used on datasets where there is a large imbalance between the positive and negative class. We highlight this in Tab. S1 and Tab. S2 for both VLM classification and object detection. In Tab. S2 in particular, the limitations of the AuROC metric are highlighted – VLM detectors with larger numbers of OSE achieve greater AuROC performance than detectors with lower OSE counts. Due to these shortcomings of AuROC, we prefer AuPR for the discussion of experiments in the main paper.
Table S1: Comparison between AuPR and AuROC metrics for the state-of-the-art VLM classifiers when tested on ImageNet1k. Arrows indicate direction of optimal performance.
AuPR AuROC TP OSE
Classifier Uncertainty $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$
CLIP [35] Softmax 71.5 79.7 31,023 50,000
Cosine 61.3 72.2
Entropy 72.3 80.2
ALIGN [21] Softmax 74.7 81.0 32,618 50,000
Cosine 62.1 72.4
Entropy 75.0 80.9
ImageBind [16] Softmax 79.1 82.8 38,405 50,000
Cosine 73.9 79.2
Entropy 80.1 84.3
@@
ViLD (R152)[17] Softmax “background" 60.1 97.0 16,011 1,485,600
Cosine “background" 33.1 96.7
Entropy “background" 62.2 97.3
@@
We test the VLM classifiers on three popular domain-specific classification datasets:
@@
In Fig. S1 and Fig. S2, we reproduce our experiments on the relationship between closed-set and open-set performance and the impact of query set size (from Sec. 5.3 and Sec. 5.4 respectively).
@@
Figure S2: Results for the domain-specific datasets when investigating the influence of query set size on closed-set and open-set performance.
S5 Investigating Individual Random Embedding Efficacy
As mentioned in Sec. 6, some negative embeddings are more effective than others at capturing open-set errors. In Fig. S3, we plot the closed-set versus open-set trade-off for $2500$ random embeddings when used with the VLM classifiers. The best performing embedding should capture many open-set images without capturing any closed-set images. We use green shading to show embeddings that capture more open-set images than closed-set images, and red shading to show the opposite – nearly all embeddings, despite being randomly selected, are in the green shading.
Figure S3: Trade-off of individual negative embeddings for capturing closed-set images incorrectly versus capturing open-set images correctly. A point in the red shading indicates the embedding is more likely to capture a closed-set image than an open-set image.
S6 Uncertainty Histograms
In Fig. S5 and Fig. S5, we show histograms visualising prediction uncertainty for the VLM classifiers and detectors respectively. We compare the uncertainty from all correct predictions when completing the closed-set test protocol with the uncertainty from all open-set errors when completing the open-set test protocol (as outlined in Sec. 4.1). In particular, the histograms highlight the different uncertainty behaviours of VLM classifiers and detectors – while the VLM classifiers are more prone to overconfident open-set errors (i.e. open-set errors with very high softmax score), the VLM detectors are more prone to underconfident correct closed-set predictions (i.e. correct predictions with low softmax score).
Figure S4: The VLM classifiers suffer from overconfident open-set errors, i.e. open-set errors with very high softmax score.
Figure S4: The VLM classifiers suffer from overconfident open-set errors, i.e. open-set errors with very high softmax score.
Figure S5: The VLM object detectors are prone to underconfident correct closed-set predictions, i.e. correct predictions with low softmax score. Note that the histograms show normalised counts for ease of visualisation, as there is large imbalance in the magnitude of the correct closed-set and open-set error sets.
S7 Precision-Recall Curves
In Fig. S6 and Fig. S7, we show the precision-recall curves for the VLM classifiers and detectors respectively. We show the precision-recall curves for the standard configuration of each VLM using different uncertainty baselines (left), the influence of different numbers of negative random words (middle) and the influence of different numbers of negative random embeddings (right).
Figure S6: Precision-recall curves for the tested VLM image classifiers.
Figure S7: Precision-recall curves for the tested VLM object detectors.
S8 Qualitative Examples
@@
In Fig. S8, we show a selection of images containing open-set errors from each of the VLM object detectors. We threshold predictions from each VLM object detector at its individual $95\%$ precision confidence threshold – this means that the open-set errors shown are the most confident and pervasive open-set errors. In some cases the open-set objects are semantically close to the predicted class (e.g. cars predicted as trucks), although in some cases this is not the case (e.g. rooster predicted as horse, or stool predicted as wine glass).

### v2

@@ 5.3 Is a good classifier all you need for VLM Open-set Recognition?
Classification: The correlation observed by [41] appears to hold for VLM classifiers, with greater accuracy (closed-set performance) relating to greater AuPR scores (open-set performance). It is worth noting that AuPR is a summary metric and this behaviour may not hold when examining performance at different operating points on the precision-recall curves.
@@ 5.4 How Does Query Set Size Impact Performance?
Figure 5: Influence of the query set size on closed-set and open-set performance on ImageNet1k classification, showing mean and standard deviation from 10 random seeds.
In Fig. 5, we test performance of the VLM classifiers on ImageNet while progressively increasing the query set size. We select the additional classes from the WordNet hierarchy by choosing leaf nodes that are related, but not too closely related, to existing ImageNet classes (i.e. share a grandparent in the hierarchy but do not share a direct parent). Increasing the number of target classes and thus the query set size decrements both the closed-set and open-set performance of a VLM classifier. This indicates that testing VLMs with an entire vocabulary of nouns may not be a feasible approach to avoid open-set recognition.
6 Discussion
Open Challenges: The tested baseline uncertainty measures were ineffective at identifying open-set errors, indicating a need to research better alternatives. While uncertainty representations for VLMs have been proposed [18, 40], their utility for open-set recognition remains unexplored.
While the use of negative random embeddings was able to reduce open-set error, large numbers were typically required to have any effect and this was often at the cost of falsely captured true positive predictions. Interestingly, some negative embeddings were much more effective at capturing open-set errors than others (e.g. the random string “awy” tested with ImageBind, see the Supplementary Sec. S9). Identifying the characteristics of these embeddings, learning them through prompt-tuning [12, 11, 8, 58, 21] or introducing auto-labelling techniques [50] are promising directions.
Conclusions: VLMs are increasingly applied in diverse real-world applications, from robotics to medical image analysis. We hope that by highlighting their weaknesses, we encourage further research to improve these models and enhance their positive societal impact. We have clearly demonstrated that VLMs for open-vocabulary image classification and object detection introduce closed-set assumptions by relying on a finite query set. We thus refuted the assumption that vision-language models (VLMs) are inherently open-set models due to their training on extensive internet-scale datasets. While the investigated baseline approaches to identify open-set errors showed promise, we found none of them sufficiently robust and reliable for safety-critical applications. Our new benchmark and evaluation protocol is designed to foster and support this research. Understanding VLMs’ limitations ultimately reduces risks, improves transparency, and ensures safer deployment in critical applications.
@@ 6 Discussion
We have made our testing and evaluation repository publicly available at github.com/dimitymiller/openset_vlms. We use the COCO API to identify True Positive (TP) predictions in object detection (see Sec. 4.1) – this is the established process for identifying TPs used during mAP calculation [24].
@@
SigLIP [53] HuggingFace Transformers google/siglip-base-patch16-224
LanguageBind [59] LanguageBind Repository LanguageBind_Image
@@
The area under the ROC curve (AuROC) is a popular metric in the traditional open-set recognition literature [7, 29, 31, 41, 54]. However a key limitation of ROC curves is that they are not affected by the size of the negative class [37] – this can make them less informative and even misleading when used on datasets where there is a large imbalance between the positive and negative class. We highlight this in Tab. S1 for VLM object detection – VLM detectors with larger numbers of OSE achieve greater AuROC performance than detectors with lower OSE counts. Due to these shortcomings of AuROC, we prefer AuPR for the discussion of experiments in the main paper.
@@
ViLD (R152)[15] Softmax “background” 60.1 97.0 16,011 1,485,600
Cosine “background” 33.1 96.7
Entropy “background” 62.2 97.3
@@
We test three VLM classifiers – CLIP [34], ALIGN [19] and ImageBind [14] – on an additional three popular domain-specific classification datasets:
@@
In Fig. S1, we reproduce our experiments on the relationship between closed-set and open-set performance from Sec. 5.3.
@@
S5 Uncertainty for Open-set and Closed-set Errors
We compare uncertainty performance for identifying errors in open-set recognition (i.e. open-set errors) versus standard closed-set recognition (i.e. closed-set misclassifications). We report the softmax and sigmoid uncertainty baselines for the VLM classifiers and object detectors.
Tab. S2 shows the results for VLM classification, where the performance of uncertainty clearly degrades for open-set error compared to closed-set error for all six classifiers. In particular, thresholding for high precision of predictions (i.e. low error rates) shows substantial performance differences, indicating there may be a greater portion of overconfident open-set errors than closed-set errors. Interestingly, SigLIP with the sigmoid activation and uncertainty appears to show the least discrepancy between closed-set error and open-set error performance; In fact, the AuROC performance for open-set errors is better than for closed-set error. This suggests that sigmoid uncertainty may not be a powerful indicator for closed-set error identification.
Table S2: Difference in uncertainty (Softmax or Sigmoid) performance between open-set error and closed-set error identification for the VLM classifiers tested on ImageNet1k.
AuROC AuPR P@95R R@95P TP Error Count
Classifier Error $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$
CLIP [34] Open-set 79.7 71.5 46.2 3.4 31,023 50,000
(Softmax) Closed-set 80.7 87.4 68.5 30.0 31,023 18,977
ALIGN [19] Open-set 81.0 74.7 48.1 8.3 32,618 50,000
(Softmax) Closed-set 81.0 89.2 70.9 35.7 32,618 17,382
CoCa [47] Open-set 79.0 71.7 46.0 3.4 31,725 50,000
(Softmax) Closed-set 81.8 88.8 70.3 33.1 31,725 18,275
ImageBind [14] Open-set 82.8 79.1 52.8 5.0 38,405 50,000
(Softmax) Closed-set 83.4 94.1 82.3 57.1 38,405 11,595
SigLIP [53] Open-set 81.0 76.2 52.1 2.2 37,851 50,000
(Sigmoid) Closed-set 70.2 87.1 78.2 5.1 37,851 12,149
LanguageBind [59] Open-set 83.9 80.7 54.7 10.5 39,243 50,000
(Softmax) Closed-set 84.2 94.8 84.1 63.3 39,243 10,757
When performing the closed-set test with the VLM object detectors, false positives can be categorised as either closed-set error or open-set error. False positives that are caused by “background” detections – detections that do not overlap a labelled object from the dataset class list – are actually open-set errors under our proposed definition. False positives that are caused by misclassification – detections that overlap a labelled object from the dataset class list but classify incorrectly – are closed-set errors. For this experiment, we identify closed-set errors as false positives from the closed-set test that overlap a labelled object with an IoU greater than $0.1$ but misclassify the object class.
Tab. S3 shows the results for VLM detection. Similar to the classification results, the performance of uncertainty clearly degrades for open-set error compared to closed-set error for all seven object detectors. For some detectors, the performance difference is substantial (e.g. the AuPR of CORA decrements by 10.2% between closed-set and open-set performance), whereas others are less notable (e.g. OV-DETR). For all detectors, one of the primary degradations from closed-set to open-set is the number of errors, with between 1.8×1.8$\times$1.8 × to 11.7×11.7$\times$11.7 × more open-set than closed-set misclassifications depending on the detector.
Table S3: Difference in uncertainty performance between open-set error and closed-set error identification for the VLM detectors tested on COCO.
AuPR P@95R R@95P TP Error Count
Detector Error $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$
OVR-CNN (R50) [51] Open-set 75.4 13.0 42.9 13,544 190,146
(Softmax) Closed-set 81.8 25.7 50.9 13,544 68,708
ViLD (R152)[15] Open-set 60.1 7.5 15.4 16,011 1,485,600
(Softmax) Closed-set 75.6 28.0 29.7 16,011 126,774
OV-DETR (R50) [49] Open-set 75.8 5.9 47.3 15,818 1,485,600
(Sigmoid) Closed-set 78.1 9.0 48.2 15,818 461,950
RegionCLIP (R50)[55] Open-set 73.1 9.9 39.4 15,741 493,940
(Softmax) Closed-set 81.6 24.1 53.4 15,741 120,483
Detic (R50)[56] Open-set 72.6 8.1 42.2 14,670 495,200
(Sigmoid) Closed-set 80.4 20.1 52.4 14,670 154,824
VLDet (R50)[22] Open-set 79.8 21.9 48.5 14,751 112,923
(Sigmoid) Closed-set 82.8 29.2 52.0 14,751 64,247
CORA (R50)[45] Open-set 67.2 8.1 24.7 13,763 495,200
(Sigmoid) Closed-set 77.4 20.0 45.8 13,763 101,015
S6 Qualitative Examples
@@
Below we show a selection of images containing open-set errors from each of the VLM object detectors. We threshold predictions from each detector at its individual $95\%$ precision confidence threshold to show the most confident and pervasive open-set errors.
(a) OVR-CNN [51].
(b) ViLD [15].
(c) OV-DETR [49].
(a) RegionCLIP [55].
(b) Detic [56]
(c) VLDet [22].
(d) CORA [45].
S7 Uncertainty Histograms
In Fig. S5 and Fig. S5, we show histograms visualising prediction uncertainty for the VLM classifiers and detectors respectively. We compare the uncertainty from all correct closed-set predictions with the uncertainty from all open-set errors. In particular, the histograms highlight the different uncertainty behaviours of VLM classifiers and detectors – while the VLM classifiers are more prone to overconfident open-set errors (i.e. open-set errors with very high softmax score), the VLM detectors are more prone to underconfident correct closed-set predictions (i.e. correct predictions with low softmax score).
Figure S4: The VLM classifiers suffer from overconfident open-set errors, i.e. open-set errors with very high softmax scores.
Figure S5: The VLM object detectors are prone to underconfident correct closed-set predictions, i.e. correct predictions with low softmax score. The histograms show normalised counts, as there is large imbalance in the magnitude of the correct closed-set and open-set error sets.
S8 Negative Embeddings and Uncertainty Interactions
For some VLM classifiers, the negative queries only capture and remove “easy” open-set errors, i.e. open-set errors that already have high uncertainty. See Fig. S6, which shows the predictions that are removed when introducing $500$ random words as negative queries. In contrast to CLIP and ALIGN, ImageBind is able to remove previously overconfident open-set errors with these negative queries, resulting in up to $5.5\%$ increase in AuPR.
Figure S6: Using negative queries (random 500 words) removes open-set error with different uncertainties depending on the VLM classifier.
S9 Investigating Individual Random Embedding Efficacy
Some negative embeddings are more effective for capturing open-set errors. Below, we plot the closed-set versus open-set trade-off for $2500$ random embeddings when used with the VLM classifiers. Ideally, an embedding should capture many open-set images without capturing any closed-set images. Green shading shows embeddings that capture more open-set than closed-set images, and red shading shows the opposite.
Figure S8: Trade-off of individual negative embeddings for capturing closed-set images incorrectly versus capturing open-set images correctly.